# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v6

@@ Abstract
Satellite Earth-observation (EO) time series in the optical and microwave ranges of the electromagnetic spectrum are often irregular due to orbital patterns and cloud obstruction. Compositing addresses these issues but loses information with respect to vegetation phenology, which is critical for many downstream tasks. Instead, we present TESSERA, a pixel-wise foundation model for multi-modal (Sentinel-1/2) EO time series that learns robust, label-efficient embeddings. During model training, TESSERA uses Barlow Twins and sparse random temporal sampling to enforce invariance to the selection of valid observations. We employ two key regularizers: global shuffling to decorrelate spatial neighborhoods and mix-based regulation to improve invariance under extreme sparsity. We find that for diverse classification, segmentation, and regression tasks, TESSERA embeddings deliver state-of-the-art accuracy with high label efficiency, often requiring only a small task head and minimal computation. To democratize access, adhere to FAIR principles, and simplify use, we release global, annual, 10m, pixel-wise int8 embeddings together with open weights/code and lightweight adaptation heads, thus providing practical tooling for large-scale retrieval and inference at planetary scale. The model training/inference code, downstream task code, and pre-generated embeddings can be accessed at https://github.com/ucam-eo.
@@ 1 Introduction
Third, unlike nearly all prior work, open-source global, annual, pixel-wise int8 embeddings for each 10m $\times$ 10m location on land, along with open-source model weights, code, and the GeoTessera Python library***https://github.com/ucam-eo/geotessera to efficiently retrieve embeddings from our data store for large-scale inference. TESSERA embeddings can be paired with lightweight adaptation heads (small MLPs or shallow convolutional models) to achieve high performance with minimal computation, enabling rapid prototyping and straightforward deployment. TESSERA thus emphasizes usability through the FAIR principles (Findable, Accessible, Interoperable, and Reusable).
@@ 2.1 Foundation Models for Earth Observation
Despite these advances, most RSFMs are trained on patch-based inputs and implicitly assume that the data have been preprocessed (e.g., aggressive cloud filtering or temporal averaging). Such assumptions simplify pretraining but discard partially cloudy images that still carry useful temporal signals, biasing representations toward idealized, cloud-free conditions. Moreover, strong performance typically requires fine-tuning the backbone for each downstream task, which can be compute- and label-intensive for many EO users.
@@ 2.2 Embedding-Based Foundation Models
A complementary line of work delivers embeddings-as-data: precomputed, generic, task-agnostic features that can be easily adapted with lightweight heads (shown in Figure 2 c). Representative models include Presto [88], which learns pixel-level temporal representations, and Google Deepmind’s AlphaEarth [11], which provides precomputed embeddings as a ready-for-analysis product. This paradigm reduces computational costs and democratizes access to high performance; however, existing approaches are often not fully open-sourced, lack pixel-level granularity at a global scale, or do not explicitly account for irregular temporal sampling.
Our work follows the embedding-based paradigm while addressing these gaps. TESSERA introduces temporal sampling invariance to handle cloud- and orbital pattern induced irregularity and combines redundancy reduction with global shuffling and mix-based regulation for robustness. In contrast to prior embedding releases, we provide FAIR global, annual, 10 m pixel-wise embeddings, together with open weights, code, and GeoTessera python library, enabling analysis-ready use without backbone fine-tuning.
@@ 3.4 Inference
Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10m pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10m resolution TESSERA embedding maps. These maps are designed as readily usable multidimensional geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing, running the TESSERA model itself (see supplementary materials). Inference is performed on 96 Intel Data Center GPU Max 1550s (128 GB GPU RAM each), processing $\sim$5 PB of raw Sentinel data per year from Microsoft Planetary Computer. Each year yields $\sim$300 TB of int8-quantized embeddings, with 2024 released first, followed by 2017–2023. We release embeddings through the pip-installable GeoTessera Python library, enabling easy access and integration into downstream workflows.
@@ 5 Discussion & Ablation Studies
Is fine-tuning ec weights necessary?
@@
Can a small spatial head recover spatial context?
Yes. Although pretraining is pixel-wise (with no explicit spatial structure), adding a very small spatial head (e.g., a small UNet) consistently improves segmentation and regression tasks, with larger gains at higher label budgets. This suggests that spatial cues can be injected downstream efficiently without complicating pretraining or creating purely pixel-based embeddings.
@@ A.3 Pre-training Details
As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training. The global batch size was 32,768. Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary fig. 10.
A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset. Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch. To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2 TB of initial d-pixel data. This pipeline is conceptually illustrated in Fig. 11a.
The impact of this approach on training stability is empirically demonstrated in Fig. 11b. Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
@@ A.4 Global Embedding Map Generation
A.5 Scaling with data and network size
To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task. Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig. 7).
@@ A.5 Scaling with data and network size
In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning. As detailed in Section A.2, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers. Our parameter sweep focused on this computationally intensive hidden block. We varied the number of these hidden layers (referred to as ’depth’ in Fig. 7a) and their width (number of neurons per layer), with results shown in Fig. 7a. As discussed by Zbontar et al. [113], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task. The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide). However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 16,384 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure. For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig. 7b. The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget. This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks. The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
A.5.1 Ablation study details
@@
A.5.2 Justification for the Number of Sampled Timesteps ($L$)
@@
Furthermore, we observed that downstream task performance is largely insensitive to the number of Sentinel-1 observations sampled (as illustrated in Figure LABEL:fig:ablation_s1_timesteps), likely due to the higher temporal density and all-weather capabilities of SAR. Therefore, for the sake of architectural consistency and to simplify the multi-modal training pipeline, we also set the number of sampled timesteps for Sentinel-1 to $L=40$.
A.6 Downstream Task Application Methodology
A.6.1 TESSERA
@@ A.6.3 Other RSFMs
A.7 Downstream Task: Austrian Crop Classification
@@
Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 14.
@@
A.8 Downstream Task: Canopy height estimation
@@
Figure 16: Comparison of canopy height estimation performance across four models. Predicted canopy height (y-axis) is compared against airborne LiDAR reference measurements (x-axis) in Danum Valley, Borneo. The figure shows scatter-density plots for four model variants trained using different input representations: TESSERA, Presto, AlphaEarth, and the DINOv2 Global CHM product (renamed from the original Meta Global CHM). All models were evaluated using the same LiDAR reference dataset for consistency. TESSERA achieves the highest coefficient of determination and lowest RMSE, demonstrating superior performance in structurally complex tropical canopies. In contrast, the DINOv2 Global CHM product exhibits strong underestimation, large negative bias, and a near-zero or negative correlation, reflecting substantial saturation at lower canopy heights. A 1:1 reference line (blue) is included in each subplot for comparison.
@@
A.9 Downstream Task: Above-ground biomass estimation
@@
Following the evaluation protocol in Ref. [63], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch. For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20). The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4). Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the AlphaEarth embeddings.
@@
A.10 Interactive Habitat Mapping Tool
@@
A.11 Pseudo Code

### v7

@@ Abstract
Satellite Earth-observation (EO) time series in the optical and microwave ranges of the electromagnetic spectrum are often irregular due to orbital patterns and cloud obstruction. Compositing addresses these issues but loses information with respect to vegetation phenology, which is critical for many downstream tasks. Instead, we present TESSERA, a pixel-wise foundation model for multi-modal (Sentinel-1/2) EO time series that learns robust, label-efficient embeddings. During model training, TESSERA uses Barlow Twins and sparse random temporal sampling to enforce invariance to the selection of valid observations. We employ two key regularizers: global shuffling to decorrelate spatial neighborhoods and mix-based regulation to improve invariance under extreme sparsity. We find that for diverse classification, segmentation, and regression tasks, TESSERA embeddings deliver state-of-the-art accuracy with high label efficiency, often requiring only a small task head and minimal computation. To democratize access, adhere to FAIR principles, and simplify use, we release global, annual, 10m, pixel-wise int8 embeddings together with open weights/code and lightweight adaptation heads, thus providing practical tooling for large-scale retrieval and inference at planetary scale. All code and data are available at https://github.com/ucam-eo/tessera.
@@ 1 Introduction
Third, unlike nearly all prior work, open-source global, annual, pixel-wise int8 embeddings for each 10m $\times$ 10m location on land, along with open-source model weights, code, and the GeoTessera Python library [58] to efficiently retrieve embeddings from our data store for large-scale inference. TESSERA embeddings can be paired with lightweight adaptation heads (small MLPs or shallow convolutional models) to achieve high performance with minimal computation, enabling rapid prototyping and straightforward deployment. TESSERA thus emphasizes usability through the FAIR principles (Findable, Accessible, Interoperable, and Reusable).
@@ 2.1 Foundation Models for Earth Observation
Despite these advances, most RSFMs are trained on patch-based input and implicitly assume that the data has been preprocessed (e.g., aggressive cloud filtering or temporal averaging). Such assumptions simplify pretraining, but discard partially cloudy images that still carry useful temporal signals, biasing representations toward idealized, cloud-free conditions. Moreover, strong performance typically requires fine-tuning the backbone for each downstream task, which can be compute- and label-intensive for EO users.
@@ 2.2 Embedding-Based Foundation Models
A complementary line of work delivers embeddings-as-data: precomputed, generic, task-agnostic features that can be easily adapted with lightweight heads (shown in Figure 2 c). Representative models include Presto [91], which learns pixel-level temporal representations, and Google’s AlphaEarth [13], which provides precomputed embeddings as a ready-for-analysis product. This paradigm reduces computational costs and democratizes access to high performance; however, existing approaches are often not fully open-sourced, lack pixel-level granularity at a global scale, or do not explicitly account for irregular temporal sampling.
Our work follows the embedding-based paradigm while addressing these gaps. TESSERA introduces temporal sampling invariance to handle cloud- and orbital pattern induced irregularity and combines redundancy reduction with global shuffling and mix-based regulation for robustness. In contrast to prior embedding releases, we provide FAIR global, annual, 10 m pixel-wise embeddings, together with open weights, code, and the GeoTessera library [58] enabling analysis-ready use without backbone fine-tuning.
@@ 3.4 Inference
Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10m pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10m resolution TESSERA embedding maps. These maps are designed as readily usable multidimensional geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing, running the TESSERA model itself (see supplementary materials). We release embeddings through the pip-installable GeoTessera Python library, enabling easy access and integration into downstream workflows.
@@ 5 Discussion & Ablation Studies
Is fine-tuning weights necessary?
@@
Can a spatial head recover spatial context?
Yes. Although pretraining is pixel-wise (with no explicit spatial structure), adding a spatial head (e.g., a UNet) consistently improves segmentation and regression tasks, with larger gains at higher label budgets. This suggests that spatial cues can be injected downstream efficiently without complicating pretraining or creating purely pixel-based embeddings.
@@ 6 Conclusion
7 Acknowledgments
We thank UKRI (MR/Z505456/1), DAWN, AMD, Vultr, the STFC Durham DiRAC HPC Facility (ST/P002293/1, ST/R002371/1, ST/S002502/1, ST/R000832/1) and the Microsoft AI For Good Lab for resources. Donations from Tarides, Jane Street, Google, Dr. Robert Sansom and John Bernstein also financially supported this research.
@@ A.3 Pre-training Details
As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training. The global batch size was 32,768. Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary  Figure 10.
A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset. Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch. To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2 TB of initial d-pixel data. This pipeline is conceptually illustrated in Figure 11 a.
The impact of this approach on training stability is empirically demonstrated in  Figure 11 b. Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
@@ A.4 Global Embedding Map Generation
A.5 Spatial Context: Pixel-wise vs. Patch-based Approaches
In this section, we elaborate on the rationale behind TESSERA’s pixel-wise design compared to traditional patch-based architectures. By maintaining pixel-wise independence, TESSERA avoids embedding fixed spatial priors, which preserves the integrity of raw spectral signatures and ensures downstream performance is driven by fundamental signal characteristics rather than pre-determined structural assumptions.
Introducing spatial context via fixed patches (Spatial variant) introduces significant vulnerabilities, particularly in regions with high cloud frequency such as Borneo. As the quantitative comparisons demonstrate (Figure 12a), introducing patch-based spatial priors can negatively impact the robustness of the representations. Furthermore, filtering for strictly clear spatial patches results in an exponential decrease in data availability, leading to severe embedding artifacts. This vulnerability is clearly visible in the embedding visualizations of the Borneo region (Figure 12b), where imposing a spatial patch requirement drastically reduces valid data and creates massive spatial voids. By extracting features at the pixel level, TESSERA allows researchers to incorporate domain-specific spatial constraints post hoc without compromising the foundational embeddings.
Figure 12: Spatial Context Ablation. (a) Quantitative comparison detailing performance differences when introducing patch-based spatial priors. (b) Visualization of the generated embeddings in the Borneo region.
A.6 Geographic Generalization
To demonstrate TESSERA’s ability to generalize to diverse phenological regions and smallholder agricultural settings, we evaluate our model on the CropHarvest Togo dataset [90] and the Ethiopia Crop classification dataset  [10]. Despite using only a continuous DoY sinusoidal encoding as a weak temporal prior, TESSERA correctly handles local seasonality driven primarily by spectral trends.
Table 5 compares TESSERA against other baseline foundation models. TESSERA significantly outperforms CROMA and AlphaEarth on both benchmarks, demonstrating superior generalization in challenging zero-shot and low-data transfer scenarios characteristic of Sub-Saharan Africa.
Table 5: Downstream Performance on Sub-Saharan African Benchmarks. Evaluation using F1 scores (%) for the Togo (CropHarvest) and Accuary (%) for the Ethiopia Crop dataset. TESSERA demonstrates superior adaptability to local seasonalities compared to existing models.
Model Togo (F1 %) Ethiopia (Acc %)
CROMA 79.15 /
AlphaEarth / $\sim$41
TESSERA (Ours) 83.01 44.92
A.7 Scaling with data and network size
To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task. Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Figure 7).
@@ A.7 Scaling with data and network size
In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning. As detailed in Section A.2, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers. Our parameter sweep focused on this computationally intensive hidden block. We varied the number of these hidden layers (referred to as ’depth’ in  Figure 7 a) and their width (number of neurons per layer), with results shown in  Figure 7 a. As discussed by Zbontar et al. [116], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task. The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide). However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 16,384 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure. For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in  Figure 7 b. The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget. This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks. The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
A.7.1 Ablation study details
@@
A.7.2 Justification for the Number of Sampled Timesteps ($L$)
@@
Furthermore, we observed that downstream task performance is largely insensitive to the number of Sentinel-1 observations sampled, likely due to the higher temporal density and all-weather capabilities of SAR. Therefore, for the sake of architectural consistency and to simplify the multi-modal training pipeline, we also set the number of sampled timesteps for Sentinel-1 to $L=40$.
A.8 Downstream Task Application Methodology
A.8.1 TESSERA
@@ A.8.3 Other RSFMs
A.9 Downstream Task: Austrian Crop Classification
@@
Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary  Figure 15.
@@
A.10 Downstream Task: Canopy height estimation
@@
Figure 17: Comparison of canopy height estimation performance across four models. Predicted canopy height (y-axis) is compared against airborne LiDAR reference measurements (x-axis) in Danum Valley, Borneo. The figure shows scatter-density plots for four model variants trained using different input representations: TESSERA, Presto, AlphaEarth, and the DINOv2 Global CHM product. All models were evaluated using the same LiDAR reference dataset for consistency. TESSERA achieves the highest coefficient of determination and lowest RMSE, demonstrating superior performance in structurally complex tropical canopies. In contrast, the DINOv2 Global CHM product exhibits strong underestimation, large negative bias, and a near-zero or negative correlation, reflecting substantial saturation at lower canopy heights. A 1:1 reference line (blue) is included in each subplot for comparison.
@@
A.11 Downstream Task: Above-ground biomass estimation
@@
Following the evaluation protocol in Ref. [65], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch. For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20). The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. Figure 5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4). Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the AlphaEarth embeddings.
@@
A.12 Interactive Habitat Mapping Tool
@@
A.13 Pseudo Code
This section provides the pseudocode for mixup regularization and global shuffling.